# Blood Flow in Elastic Vessel Using PINNH. Zhang, R. Chan and X.C. Tai

### v1

@@ 3.2 Loss Function
Figure 3: The illustration for the Network Architecture. Three networks are used. Displacement network $N_{d}$ is for solving the displacement vector $\boldsymbol{d}$ for both the fluid and the solid domain. Velocity network $N_{u}$ is to compute the fluid velocity $\boldsymbol{u}$ and pressure network $N_{p}$ is for the fluid pressure $p$. Note that the green line indicates data flow where gradient calculation needs to be done, while the black line denotes that no gradient needs to be delivered. The dotted black line means that the computed results need to be given in the loss function calculation. But it will not enable a back-propogation which might change the network producing it. For example, when solving for the fluid problem in the alternative optimization, the loss function (24) matters with $\boldsymbol{d}$. However, the gradient will not be computed with respect to $\boldsymbol{d}$ and the parameters of $N_{d}$, which is explained in more detail in Section.3.4.
@@
$\eta^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d})=N_{d}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d})\quad\text{ for }\boldsymbol{x}_{0}\in\Omega_{t};t\in[0,T].$ (18)
@@
$\boldsymbol{d}^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d})=\eta^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d})\boldsymbol{e}_{r}\quad\text{ for}\boldsymbol{x}_{0}\in\Omega_{t};t\in[0,T].$ (19)
@@
$\boldsymbol{x}_{t}^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d})=\boldsymbol{x}_{0}+\boldsymbol{d}^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d})\quad\text{ for }\boldsymbol{x}_{0}\in\Omega_{t};t\in[0,T].$ (20)
@@
$\boldsymbol{u}^{*}(\boldsymbol{x}_{t}^{*},t^{\prime};{\boldsymbol{\theta}}_{u},\boldsymbol{\theta}_{d})&=N_{u}(\boldsymbol{x}_{t}^{*},t^{\prime};{\boldsymbol{\theta}}_{u})=N_{u}(\boldsymbol{x}_{0}+\boldsymbol{d}^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d}),t^{\prime};{\boldsymbol{\theta}}_{u})\\ P^{*}(\boldsymbol{x}_{t}^{*},t^{\prime};{\boldsymbol{\theta}}_{p},\boldsymbol{\theta}_{d})&=N_{p}(\boldsymbol{x}_{t}^{*},t^{\prime};{\boldsymbol{\theta}}_{p})=N_{p}(\boldsymbol{x}_{0}+\boldsymbol{d}^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d}),t^{\prime};{\boldsymbol{\theta}}_{p})\end{aligned}\quad\text{ for }\boldsymbol{x}_{0}\in\Omega_{t}^{\text{f}};t\in[0,T].$ (21)
@@
$\mathcal{L}_{\text{ns}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})$ $=||\frac{\partial}{\partial t^{\prime}}\boldsymbol{u^{*}}+(\boldsymbol{u}^{*}\cdot\nabla)\boldsymbol{u}^{*}+\nabla p^{*}-2\nabla\cdot(\mu\boldsymbol{D}(\boldsymbol{u}^{*}))||^{2}_{L^{2}(\Omega_{t}^{\text{f}};[0,T])},$ (22)
@@
$\mathcal{L}_{\text{bdr}}^{\text{f}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})$ $=||F^{\text{f}}_{\text{bdr}}(\boldsymbol{u}^{*},P^{*})||^{2}_{L^{2}(\Gamma_{t};[0,T])},$
@@
Note that the $L^{2}$ norm is employed here and will be redefined by a discrete approximation in the next subsection.
@@
$\mathcal{L}_{\text{sc}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})$ $=||\frac{\partial^{2}\eta^{*}}{\partial t^{2}}+b\eta^{*}-H||^{2}_{L^{2}(\Gamma_{0}^{\text{w}};[0,T])},$ (23)
@@
where $\eta^{*}(\boldsymbol{x}_{0})$ is the projection of $\boldsymbol{d}^{*}(\boldsymbol{x}_{0})$ onto $\boldsymbol{e}_{r}$ for $\boldsymbol{x}_{0}\in\Gamma_{0}^{\text{w}}$ as discussed in Sec.2.
@@
$\mathcal{L}_{\text{fluid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})=\alpha_{\text{ns}}\mathcal{L}_{\text{ns}}+\alpha_{\text{bdr}}^{\text{f}}\mathcal{L}_{\text{bdr}}^{\text{f}}+\alpha_{\text{init}}^{\text{f}}\mathcal{L}_{\text{init}}^{\text{f}},$ (24)
$\mathcal{L}_{\text{solid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})=\alpha_{\text{sc}}\mathcal{L}_{\text{sc}}+\alpha_{\text{he}}\mathcal{L}_{\text{he}}+\alpha_{\text{bdr}}^{\text{s}}\mathcal{L}_{\text{bdr}}^{\text{s}}+\alpha_{\text{init}}^{\text{s}}\mathcal{L}_{\text{init}}^{\text{s}}.$ (25)
@@
$\min_{{\boldsymbol{\theta}}_{d},{\boldsymbol{\theta}_{f}}}\mathcal{L}_{\text{solid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}}),$ (26)
@@
$\min_{{\boldsymbol{\theta}_{f}},\boldsymbol{\theta}_{d}}\mathcal{L}_{\text{fluid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}}).$ (27)
Our method to solve the above FSI problem is to solve the fluid problem and the solid problem iteratively. In particular, the solution of the fluid problem is obtained by imposing the displacement obtained from solving the solid problem at the interface as boundary conditions. The result for the solid problem is solved by imposing $H$ in the stress continuity (7) on the interface. Correspondingly, in our method, which needs to minimize the two losses, we take an alternative optimization strategy to decouple the FSI problem into two sub-problems. The optimization scheme will be given more specifically in 3.4.
3.3 Mini-batch gradient descent algorithm
$\mathcal{S}({\Omega_{0}^{\text{f}}};t=0)=\{(\boldsymbol{x}^{(i)}_{0},0)\quad|\quad\text{ where }\boldsymbol{x}^{(i)}_{0}\in\mathcal{S}({\Omega_{0}^{\text{f}}})\}$ (28)
$\mathcal{S}({\Omega_{t}^{\text{f}}};[0,T])=\{(\boldsymbol{x}^{(i)}_{t^{(i)}},t^{(i)})\quad|\quad$ $\boldsymbol{x}^{(i)}_{t^{(i)}}:=\boldsymbol{x}^{(i)}_{0}+\boldsymbol{d}^{*}(\boldsymbol{x}^{(i)}_{0},t^{(i)};\boldsymbol{\theta}_{d})$ (29)
[…]
$||f||^{2}_{\mathcal{S}({\Omega_{t}^{\text{f}}};[0,T])}:=\frac{1}{M}\sum\limits_{(\boldsymbol{x},t)\in\mathcal{S}({\Omega_{t}^{\text{f}}};[0,T])}|f(\boldsymbol{x},t)|^{2}$ (30)
$||\cdot||^{2}_{L^{2}(\cdot;\cdot)}:=||\cdot||^{2}_{\mathcal{S}(\cdot;\cdot)}$ (31)
To decrease the value of the losses, gradient descent methods are applied to update ${\boldsymbol{\theta}_{f}}$ in and ${\boldsymbol{\theta}_{f}}$ by
${\boldsymbol{\theta}_{f}}\leftarrow{\boldsymbol{\theta}_{f}}-\tau_{f}\nabla_{{\boldsymbol{\theta}_{f}}}\mathcal{L}_{\text{fluid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}}),$ (32)
${\boldsymbol{\theta}_{d}}\leftarrow{\boldsymbol{\theta}_{d}}-\tau_{d}\nabla_{{\boldsymbol{\theta}_{d}}}\mathcal{L}_{\text{solid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}}),$ (33)
where $\nabla_{{\boldsymbol{\theta}_{f}}}\mathcal{L}_{\text{fluid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})$ is the gradient of $\mathcal{L}_{\text{fluid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})$ with respect to ${\boldsymbol{\theta}_{f}}$, and $\nabla_{{\boldsymbol{\theta}_{d}}}\mathcal{L}_{\text{solid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})$ is the gradient of $\mathcal{L}_{\text{solid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})$ with respect to ${\boldsymbol{\theta}_{d}}$. Here, $\tau_{f}$, $\tau_{d}$ are adaptive step lengths calculated by some methods [1, 18]. In this paper, since we utilize Adam optimizer in the implementation, the step lengths are all calculated using the method in [18].
@@ 3.4 Training Scheme
Before stepping into the training scheme, we define a training unit for the fluid problem to be a collection of $m$ training epochs to make our discussion convenient as given in Algorithm 2. Within each unit for the fluid problem, using the loss function (24), $N_{u}$ is trained for the first $m_{1}=80\%\times m$ epochs, and $N_{p}$ is trained for the remaining $m_{2}=20\%\times m$ epochs. The training unit for the solid problem is defined as a group of $m$ training epochs for $N_{p}$ using (25).
To initialize, we perform zero-initialization on the parameters for the output layer of $N_{d}$ so that its output is all zeros, aligning with the boundary and initial conditions for solving displacements in solid problems for most cases. Then, with $\alpha_{\text{ns}}=0$, we optimize $N_{u}$ and $N_{p}$ for a maximum of $m_{\text{fluid}}$ units, which can be stopped when it converges before that. The unit-by-unit training results in an alternative training for $N_{u}$ and $N_{p}$ until the optimization converges (the improvement on loss is less than a control parameter $\epsilon$). This sequence of operations ensures that the networks fulfill the initial and boundary conditions.
After the initialization phase, we further train the networks $N_{u}$ and $N_{p}$ to solve the Navier-Stokes equations in the domain of the reference frame. During this stage, we initially set $\alpha_{\text{ns}}=10^{-7}$. Then, we update it, according to Section.4.1.2, by a factor of $10$ when the loss function $\mathcal{L}_{\text{fluid}}$ converges or the training units have run for $m_{\text{fluid}}$ units. The update is performed for $E$ times for one simulation problem.
The final stage of training involves coupling the fluid and solid systems. In this phase, we alternately solve the fluid and solid problems. Specifically, we run $m_{\text{solid}}$ training units to solve the solid problem for the displacement computation, followed by a maximum of $m_{\text{fluid}}$ training units for the fluid problem. Either training can be stopped in advance in one alternation if it converges. We alternate between these two problems until they converge or run for a maximum of $F$ alternations.
@@
2:Set $\alpha_{\text{ns}}\leftarrow 0$;
3:Optimize the fluid problem for $m_{\text{fluid}}$ training units at most;
4:Take $\alpha_{\text{ns}}\leftarrow 10^{-8}$;
@@
7:     Optimize the fluid problem for $m_{\text{fluid}}$ training units at most;
@@
10:     Optimize the solid problem for $m_{\text{solid}}$ training units at most;
11:     Optimize the fluid problem for $m_{\text{fluid}}$ training units at most;
@@
2:while $|L^{(k)}-L^{(k-1)}|<\epsilon$ or $k>m_{\text{fluid}}$ do
3:     Train $N_{u}$ for $m_{1}$ epochs, and then $N_{p}$ for $m_{2}$ epochs using $\mathcal{L}_{\text{fluid}}$;
4:     $L^{(k)}\leftarrow$ the minimal value for $\mathcal{L}_{\text{fluid}}$ within this training unit;
5:     $k=k+1$;
6:end while
3.5 Parallel Training
$\nabla_{{\boldsymbol{\theta}_{d}}}\mathcal{L}_{\text{solid}}(\boldsymbol{\theta}_{\text{d}};\mathcal{N})=\frac{1}{P}\sum\limits_{i=1}^{P}\nabla_{{\boldsymbol{\theta}_{\text{d}}^{(i)}}}\mathcal{L}_{\text{solid}}(\boldsymbol{\theta}_{\text{d}}^{(i)};\mathcal{N}^{(i)})$ (34)

### v4

@@ 3.2 Loss Function
Figure 2: Illustration of Network Architecture: Three networks are employed. The displacement network $N_{d}$ solves the displacement vector $\boldsymbol{d}$ for both fluid and solid domains. The velocity network $N_{u}$ computes fluid velocity $\boldsymbol{u}$, while the pressure network $N_{p}$ handles fluid pressure $p$. Green lines indicate data flow requiring gradient calculation, while black lines denote no gradient delivery. Dotted black lines indicate computed results for loss function calculation without enabling back-propagation, as explained in Section 3.3.
@@
(18) $\eta^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d})=N_{d}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d})\quad\text{ for }\boldsymbol{x}_{0}\in\Omega_{t};t\in[0,T].$
@@
(19) $\boldsymbol{d}^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d})=\eta^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d})\boldsymbol{e}_{r}\quad\text{ for}\boldsymbol{x}_{0}\in\Omega_{t};t\in[0,T].$
@@
(20) $\boldsymbol{x}_{t}^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d})=\boldsymbol{x}_{0}+\boldsymbol{d}^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d})\quad\text{ for }\boldsymbol{x}_{0}\in\Omega_{t};t\in[0,T].$
@@
(21) $\boldsymbol{u}^{*}(\boldsymbol{x}_{t}^{*},t^{\prime};{\boldsymbol{\theta}}_{u},\boldsymbol{\theta}_{d})&=N_{u}(\boldsymbol{x}_{t}^{*},t^{\prime};{\boldsymbol{\theta}}_{u})=N_{u}(\boldsymbol{x}_{0}+\boldsymbol{d}^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d}),t^{\prime};{\boldsymbol{\theta}}_{u})\\ P^{*}(\boldsymbol{x}_{t}^{*},t^{\prime};{\boldsymbol{\theta}}_{p},\boldsymbol{\theta}_{d})&=N_{p}(\boldsymbol{x}_{t}^{*},t^{\prime};{\boldsymbol{\theta}}_{p})=N_{p}(\boldsymbol{x}_{0}+\boldsymbol{d}^{*}(\boldsymbol{x}_{0},t;\boldsymbol{\theta}_{d}),t^{\prime};{\boldsymbol{\theta}}_{p})\end{aligned}\quad\text{ for }\boldsymbol{x}_{0}\in\Omega_{t}^{\text{f}};t\in[0,T].$
@@
(22) $\mathcal{L}_{\text{ns}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})$ $=||\frac{\partial}{\partial t^{\prime}}\boldsymbol{u^{*}}+(\boldsymbol{u}^{*}\cdot\nabla)\boldsymbol{u}^{*}+\nabla P^{*}-2\nabla\cdot(\mu\boldsymbol{D}(\boldsymbol{u}^{*}))||^{2}_{L^{2}(\Omega_{t}^{\text{f}};[0,T])},$
@@
$\mathcal{L}_{\text{bdr}}^{\text{f}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})$ $=||F^{\text{f}}_{\text{bdr}}(\boldsymbol{u}^{*},P^{*})||^{2}_{L^{2}(\Gamma_{t}^{\text{w}};[0,T])},$
@@
Note that the $L^{2}$ norm is employed here and will be redefined by a discrete approximation in the next subsection. Besides, since we assume the problem to be axis-symmetric, we also include the cylindrical form for the Navier-Stokes equation in Appendix.
@@
(23) $\mathcal{L}_{\text{sc}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})$ $=||\frac{\partial^{2}\eta^{*}}{\partial t^{2}}+b\eta^{*}-H||^{2}_{L^{2}(\Gamma_{0}^{\text{w}};[0,T])},$
@@
where $\eta^{*}(\boldsymbol{x}_{0})$ is the projection of $\boldsymbol{d}^{*}(\boldsymbol{x}_{0})$ onto $\boldsymbol{e}_{r}$ for $\boldsymbol{x}_{0}\in\Gamma_{0}^{\text{w}}$ as discussed in Sec.2. The cylindrical form for the harmonic extension calculation $\mathcal{L}_{\text{he}}$ is given in Appendix.
@@
(24) $\mathcal{L}_{\text{fluid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})=\alpha_{\text{ns}}\mathcal{L}_{\text{ns}}+\alpha_{\text{bdr}}^{\text{f}}\mathcal{L}_{\text{bdr}}^{\text{f}}+\alpha_{\text{init}}^{\text{f}}\mathcal{L}_{\text{init}}^{\text{f}},$
(25) $\mathcal{L}_{\text{solid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})=\alpha_{\text{sc}}\mathcal{L}_{\text{sc}}+\alpha_{\text{he}}\mathcal{L}_{\text{he}}+\alpha_{\text{bdr}}^{\text{s}}\mathcal{L}_{\text{bdr}}^{\text{s}}+\alpha_{\text{init}}^{\text{s}}\mathcal{L}_{\text{init}}^{\text{s}}.$
@@
(26) $\min_{{\boldsymbol{\theta}}_{d},{\boldsymbol{\theta}_{f}}}\mathcal{L}_{\text{solid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}}),$
@@
(27) $\min_{{\boldsymbol{\theta}_{f}},\boldsymbol{\theta}_{d}}\mathcal{L}_{\text{fluid}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}}).$
To decrease the value of the losses, gradient descent methods are applied to update ${\boldsymbol{\theta}_{f}}=\{\boldsymbol{\theta}_{u},\boldsymbol{\theta}_{p}\}$ and ${\boldsymbol{\theta}_{d}}$ using three different optimizer by
(28) $\boldsymbol{\theta}_{u}\leftarrow\boldsymbol{\theta}_{u}-\tau_{u}\nabla_{\boldsymbol{\theta}_{u}}\mathcal{L}_{\text{fluid}}(\boldsymbol{\theta}_{f},\boldsymbol{\theta}_{d}),$
(29) $\boldsymbol{\theta}_{p}\leftarrow\boldsymbol{\theta}_{p}-\tau_{p}\nabla_{\boldsymbol{\theta}_{p}}\mathcal{L}_{\text{fluid}}(\boldsymbol{\theta}_{f},\boldsymbol{\theta}_{d}),$
(30) $\boldsymbol{\theta}_{d}\leftarrow\boldsymbol{\theta}_{d}-\tau_{d}\nabla_{\boldsymbol{\theta}_{d}}\mathcal{L}_{\text{solid}}(\boldsymbol{\theta}_{f},\boldsymbol{\theta}_{d}),$
where $\nabla_{\boldsymbol{\theta}_{u}}$, $\nabla_{\boldsymbol{\theta}_{p}}$ and $\nabla_{\boldsymbol{\theta}_{d}}$ are the gradients with respect to $\boldsymbol{\theta}_{u}$, $\boldsymbol{\theta}_{p}$ and $\boldsymbol{\theta}_{d}$. Here, $\tau_{u}$, $\tau_{p}$, $\tau_{d}$ are adaptive step lengths calculated by some methods [1, 20]. In the implementation, we utilize three Adam optimizers for each of the gradient descents above, whose step lengths are calculated using the method in [20].
Our method to solve the above FSI problem is to solve the fluid problem and the solid problem iteratively. In particular, the solution of the fluid problem is obtained by imposing the displacement obtained from solving the solid problem at the interface as boundary conditions. The result for the solid problem is solved by imposing $H$ in the stress continuity (7) on the interface. Correspondingly, in our method, which needs to minimize the two losses, we take an alternative optimization strategy to decouple the FSI problem into two sub-problems. The optimization scheme will be given more specifically in 3.3.
@@ 3.3 Training Scheme
We first apply zero-initialization to the parameters of the output layer of $N_{d}$, which lets the output all zeros. This initialization allows the solid domain to start in a non-deformed state, aligned with the reference domain and provides a controlled start to the simulation process. Then, with $\alpha_{\text{ns}}=0$, we optimize $N_{u}$ and $N_{p}$ for $m_{\text{fluid}}$ epochs alternatively with an alternation every $m_{1}+m_{2}$ epochs as indicated by Algorithm 2. This sequence of operations ensures that the networks fulfill the initial and boundary conditions.
After the initialization phase, we further train the networks $N_{u}$ and $N_{p}$ to solve the Navier-Stokes equations in the domain of the reference frame. During this stage, we initially set $\alpha_{\text{ns}}=10^{-7}$. Then, we update it by a factor of $10$ when the loss function $\mathcal{L}_{\text{fluid}}$ converges or the training has run for $m_{\text{fluid}}$ epochs. The $\alpha_{\text{ns}}$ update is performed for $E$ times for one simulation problem. This strategy is specifically crafted to address difficulties in decreasing losses when dealing with multiple losses. By weighting the loss terms sequentially, we aim to simplify the optimization process and enhance convergence, allowing the optimizer to focus on more heavily weighted losses.
The final stage of training involves coupling the fluid and solid systems. In this phase, we alternately solve the fluid and solid problems. Specifically, we run $m_{\text{solid}}$ epochs to solve the solid problem for the displacement computation, followed by a maximum of $m_{\text{fluid}}$ epochs for the fluid problem. Either training can be stopped in advance in one alternation if it converges (when the error improvement is less than $\epsilon$ for $100$ epochs). We alternate between these two problems until they converge or run for a maximum of $F$ alternations. The decision to train the networks alternatively is based on the concept of multi-objective optimization, which often employs a variety of alternative techniques to find solutions efficiently. This approach allows for more stable training and better convergence by focusing on optimizing subsets of losses sequentially.
@@
2:$\alpha_{\text{ns}}\leftarrow 0$;
3:$\boldsymbol{\theta}_{u},\boldsymbol{\theta}_{p}\leftarrow$ Optimize fluid problem using Alg. 2 for $m_{\text{fluid}}$ epochs;
4:$\alpha_{\text{ns}}\leftarrow 10^{-8}$;
@@
7:     $\boldsymbol{\theta}_{u},\boldsymbol{\theta}_{p}\leftarrow$ Optimize fluid problem using Alg. 2 for $m_{\text{fluid}}$ epochs;
@@
10:     $\boldsymbol{\theta}_{d}\leftarrow$ Optimize solid problem using (30) for $m_{\text{solid}}$ epochs;
11:     $\boldsymbol{\theta}_{u},\boldsymbol{\theta}_{p}\leftarrow$ Optimize fluid problem using Alg. 2 for $m_{\text{fluid}}$ epochs;
@@
2:for $k=1:\frac{m_{\text{fluid}}}{(m1+m2)}$ do
3:     $\boldsymbol{\theta}_{u}\leftarrow$ Optimize fluid problem using (28) for $m_{1}$ epochs;
4:     $\boldsymbol{\theta}_{p}\leftarrow$ Optimize fluid problem using (29) for $m_{2}$ epochs;
5:end for